### v2

@@ Abstract
Recent advances in learned image compression (LIC) have achieved remarkable performance improvements over traditional codecs. Notably, the MLIC series—LICs equipped with multi-reference entropy models—have substantially surpassed conventional image codecs such as Versatile Video Coding (VVC) Intra. However, existing MLIC variants suffer from several limitations: performance degradation at high bitrates due to insufficient transform capacity, suboptimal entropy modeling that fails to capture global correlations in initial slices, and lack of adaptive channel importance modeling. In this paper, we propose MLICv2 and MLICv2+, enhanced successors that systematically address these limitations through improved transform design, advanced entropy modeling, and exploration of the potential of instance-specific optimization. For transform enhancement, we introduce a lightweight token mixing block inspired by the MetaFormer architecture, which effectively mitigates high-bitrate performance degradation while maintaining computational efficiency. For entropy modeling improvements, we propose hyperprior-guided global correlation prediction to extract global context even in the initial slice of latent representation, complemented by a channel reweighting module that dynamically emphasizes informative channels. We further explore enhanced positional embedding and guided selective compression strategies for superior context modeling. Additionally, we apply the Stochastic Gumbel Annealing (SGA) to demonstrate the potential for further performance improvements through input-specific optimization. Extensive experiments demonstrate that MLICv2 and MLICv2+ achieve state-of-the-art results, reducing Bjøntegaard-Delta Rate by 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% on Kodak, Tecnick, and CLIC Pro Val datasets, respectively, compared to VTM-17.0 Intra.
@@ 1 Introduction
Image compression constitutes a fundamental challenge in multimedia communication, with the objective of achieving compact representation while preserving reconstruction quality. Traditional compression methods, exemplified by VVC Intra [12], rely on block-based hybrid coding frameworks refined over decades. However, as conventional approaches near performance saturation, learned image compression (LIC) has emerged as a transformative paradigm, demonstrating substantial potential to surpass hand-crafted codecs through end-to-end optimization.
Contemporary LIC methods [9, 64, 10, 50, 51, 16, 75, 31, 30, 88, 35, 86, 68, 87, 38, 19, 71, 26, 74, 78, 53, 15, 45, 34, 33, 32] predominantly adopt autoencoder-based frameworks with rate-distortion optimization, comprising analysis transforms, quantization, synthesis transforms, and entropy models. The analysis transform maps input images to latent representations, which undergo quantization before reconstruction via synthesis transforms. Entropy models estimate quantized latent distributions to facilitate efficient entropy coding. Advanced techniques of transform architectures and entropy models have enabled several LICs [30, 34, 33, 32, 45, 40, 47, 20] to achieve performance comparable to or exceeding advanced traditional codecs like VVC Intra [12].
@@
Figure 1: BD-Rate and decoding time during inference on CLIC Professional Valid [65] with 2K resolution. The bubble size indicates the number of model parameters.
Entropy modeling plays a pivotal role in LIC performance, aiming to minimize conditional entropy through informative prior capture. While hyperprior models [10] provide side information, context models [52, 31] capture spatial dependencies between previously decoded and current elements. Building on these foundations, Jiang et al. introduced the MLIC series [34, 33], which leverages multi-reference priors through latent representation partitioning into channel-wise slices [51], capturing both inter-slice and intra-slice dependencies. MLIC employs checkerboard partitioning [31] within slices for two-pass decoding, while attention mechanisms model global dependencies from previous slices. MLIC++ [33] addressed computational complexity by replacing vanilla attention [70] with linear attention [60] and modeling global correlations across previous slices.
@@
Despite the strong overall performance of the MLIC series, several critical limitations persist and motivate our work. First, regarding transform design, MLIC models exhibit performance degradation at high bitrates, primarily attributed to insufficient transform capacity for preserving latent representation information. Second, concerning entropy modeling, global correlation estimation relies exclusively on previous slices, precluding global correlation capture in initial slices. Moreover, existing approaches treat all channels equally without considering varying informational importance, potentially limiting entropy modeling accuracy.
@@
To systematically address these limitations, we propose MLICv2, featuring enhanced transform and entropy modeling modules, and MLICv2+, which additionally incorporates iterative refinement techniques [80] to demonstrate framework potential.
@@
For transform improvement, we design an efficient token mixing block replacing residual blocks in MLIC++. Inspired by MetaFormer architecture [81], our block employs depth-wise convolution for spatial token mixing and lightweight gated interaction for point-wise operations, achieving superior high-bitrate performance with reduced complexity.
@@
For entropy modeling enhancement, we propose MEMv2, an advanced multi-reference entropy model capturing diverse dependencies. We introduce hyperprior-guided global correlation prediction for initial slices, leveraging spatial distribution similarity between hyperprior and latent representations. Our two-stage context modeling strategy performs spatial aggregation followed by adaptive channel-wise reweighting to capture feature importance variations. Additional improvements include 2D Rotary Positional Embedding (RoPE) and guided selective compression for enhanced modeling and coding efficiency.
@@
Additionally, we employ Stochastic Gumbel Annealing (SGA) [80] to iteratively refine latent representations and side information for each input. This approach demonstrates our framework’s potential for input-specific rate-distortion optimization while preserving decoder complexity, suggesting opportunities for further architectural enhancements.
@@
Our comprehensive approach yields significant improvements, with MLICv2 and MLICv2+ substantially outperforming MLIC++ as demonstrated in Figure 1. MLICv2 achieves BD-rate reductions of 16.54%, 21.61%, and 16.05% on Kodak, Tecnick, and CLIC Pro Val datasets respectively, while MLICv2+ further improves these results to 20.46%, 24.35%, and 19.14% over VTM-17.0 Intra. Our contributions are summarized as follows:
@@
  • We introduce MetaFormer-based architecture for LIC through our simple token mixing block, achieving superior high-bitrate performance with reduced computational complexity compared to previous residual blocks.
@@
  • We systematically improve entropy modeling through multiple improvements: (i) hyperprior-guided global correlation prediction addressing initial slice limitations, (ii) the first two-stage context modeling strategy with spatial aggregation and adaptive channel reweighting, and (iii) advanced positional embedding and selective compression mechanisms.
  • We demonstrate the effectiveness of our framework by applying iterative refinement techniques (SGA), showing consistent performance gains without decoder complexity increase and suggesting opportunities for further architectural enhancements.
@@
  • Extensive experiments validate our approach, with MLICv2 and MLICv2+ achieving BD-rate reductions of 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% over VTM-17.0 Intra on Kodak, Tecnick, and CLIC Pro Val datasets respectively.
@@ 2.1 Learned Image Compression
Autoencoder-based learned image compression [9, 64, 50, 16, 30, 34, 33, 51, 32] follows the transform coding paradigm, consisting of an analysis transform $g_{a}$ with parameters $\phi$, a synthesis transform $g_{s}$ with parameters $\psi$, a quantization function $\lceil\cdot\rfloor$, and an entropy model $p$ with parameters $\rho$. Given an input image $\boldsymbol{x}\sim p_{\boldsymbol{x}}$, the process is formulated as:
$\boldsymbol{y}=g_{a}(\boldsymbol{x};\phi),\hat{\boldsymbol{y}}=\lceil\boldsymbol{y}\rfloor,\hat{\boldsymbol{x}}=g_{s}(\hat{\boldsymbol{y}};\psi).$ (1)
The rate-distortion optimization is achieved through:
$\mathcal{L}=\mathbb{E}_{\boldsymbol{x}\sim p_{\boldsymbol{x}}}\left[\underbrace{-\log_{2}\int p(\hat{\boldsymbol{y}};\rho)d\hat{\boldsymbol{y}}}_{\textrm{bit-rate}}+\lambda\times\underbrace{\mathcal{D}(\hat{\boldsymbol{x}},\boldsymbol{x})}_{\textrm{distortion}}\right],$ (2)
where $\lambda$ balances the rate-distortion trade-off. During training, quantization is approximated by adding uniform noise [9] or straight-through estimation [64]. Ballé et al. [10] introduce hyperprior networks using hyper analysis $h_{a}$ and hyper synthesis $h_{s}$ to improve entropy estimation through quantized side information $\hat{\boldsymbol{z}}$ extracted from latent representations.
Transform architectures have evolved significantly, incorporating Generalized Divisive Normalization (GDN) [8], residual non-local attention [85, 15], deep residual networks [16], and transformer-based designs [87, 88, 48]. Mixed CNN-transformer [45] and Mamba [82, 25, 56] architectures have also been explored for capturing both local and non-local interactions. Entropy estimation has progressed from simple univariate Gaussian models [10] to sophisticated approaches including mean-scale Gaussian models [50], asymmetric Gaussian distributions [17], Gaussian mixture models [16], and generalized Gaussian models [83], significantly improving compression efficiency.
@@ 2.2 Conditional Entropy Models
Local context has evolved from serial approaches like PixelCNN [69, 50] to efficient parallel methods including zigzag scanning [41] and checkerboard partitioning [30]. These intra-slice local contexts are complemented by inter-slice approaches and channel-wise autoregressive coding [51] for faster decoding.
Recent works investigate long-range correlations among distant elements [55, 26, 36, 34, 33]. Qian et al. [55, 54] employ neighbor elements and bi-directional attention for global correlation estimation. Guo et al. [26] utilize L-2 distance metrics, while Kim et al. [36] transmit global information to decoders. Recently, Jiang et al. [34] propose MLIC and MLIC+, dividing global contexts into inter-slice and intra-slice components. The latent representation is partitioned into multiple slices for autoregressive conditional coding, employing checkerboard attention for intra-slice contexts and convolutional layers for inter-slice contexts. However, the quadratic complexity of vanilla attention [70] limits scalability for high-resolution images, which is addressed by MLIC++ [33] through linear attention mechanisms [60].
Figure 2: The overall architecture of MLICv2/MLICv2+. $g_{a}$ is the analysis transform, $g_{s}$ is the synthesis transform, $h_{a}$ is the hyper analysis, and $h_{s}$ is the hyper synthesis. $g_{ep}$ is the entropy parameter module. $f_{s}$ is the selective compression predictive module. $g_{lrp}$ is the latent residual prediction module [51]. AE and AD are arithmetic encoding and decoding. $\uparrow$ denotes upsampling and $\downarrow$ denotes downsampling. $\boldsymbol{x}$} is the input image and $\hat{\boldsymbol{x}}$ is the reconstructed image. $\boldsymbol{y}$ is the latent representation and $\hat{\boldsymbol{y}}$ is the quantized latent representation. $\hat{\boldsymbol{y}}^{i}$ is the $i$-th slice of $\hat{\boldsymbol{y}}$. $\boldsymbol{\mu}^{i},\boldsymbol{\sigma}^{i}$ are the mean and scale of $\hat{\boldsymbol{y}}^{i}$. $\boldsymbol{s}^{i}$ is the predicted selective coding map. $\boldsymbol{r}^{i}$ is the predicted residual. $\hat{\boldsymbol{z}}$ is the side information. $M,N$ are the channel numbers, which are $320,192$ in MLICv2/MLICv2+. “Refine” is employed in MLICv2+.
@@ 3.1 Background: MLIC++
Our proposed MLICv2 and MLICv2+ build upon the foundation established by MLIC++ [33]. MLIC++ employs a simplified version of Cheng’20’s transform [16], where attention modules are removed to reduce computational complexity. For entropy modeling, MLIC++ introduces a linear-complexity multi-reference entropy model (MEM++), which effectively captures hyperprior, intra-slice, and inter-slice dependencies.
The hyperprior information $\hat{\boldsymbol{H}}$ is derived from side information $\hat{\boldsymbol{z}}$. Following the approach of Minnen et al. [51], the latent representation $\hat{\boldsymbol{y}}$ is partitioned into $L+1$ slices: $\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\ldots,\hat{\boldsymbol{y}}^{L}$}. Each slice $\hat{\boldsymbol{y}}^{i}$ is further divided into anchor $\hat{\boldsymbol{y}}^{i}_{ac}$ and non-anchor $\hat{\boldsymbol{y}}^{i}_{na}$ components using checkerboard partitioning [31].
The inter-slice local context module $g_{\ell}^{inter}$ captures inter-slice local context $\hat{\boldsymbol{C}}^{<i}_{\ell}$ from preceding slices $\hat{\boldsymbol{y}}^{<i}$. To model inter-slice global context $\hat{\boldsymbol{C}}^{<i}_{g}$, MLIC++ employs linear attention [60] to achieve an optimal balance between performance and computational complexity. The intra-slice local context $\hat{\boldsymbol{C}}^{i}_{\ell}$ is extracted from the anchor partition $\hat{\boldsymbol{y}}^{i}_{ac}$ using overlapped window-based checkerboard attention $g_{\ell}^{intra}$.
@@
The estimated means $\boldsymbol{\mu}^{i}_{ac}$, $\boldsymbol{\mu}^{i}_{na}$ and scales $\boldsymbol{\sigma}^{i}_{ac}$, $\boldsymbol{\sigma}^{i}_{na}$ are extracted from the contextual information using the entropy parameter module $g_{ep}$. MLIC++ incorporates mixed quantization [51] and the latent residual prediction module $g_{lrp}$ [51]. In mixed quantization, zero-centered straight-through estimation (STE) [64] is applied for distortion computation, while additive uniform noise (AUN) [9] is used for entropy estimation during training. The module $g_{lrp}$ predicts quantization residuals based on decoded slices and hyperprior information:

### v3

@@ Abstract
Recent advances in learned image compression (LIC) have achieved remarkable performance improvements over traditional codecs. Notably, the MLIC series—LICs equipped with multi-reference entropy models—have substantially surpassed conventional image codecs such as Versatile Video Coding (VVC) Intra. However, existing MLIC variants suffer from several limitations: performance degradation at high bit-rates due to insufficient transform capacity, suboptimal entropy modeling that fails to capture global correlations in initial slices, and lack of adaptive channel importance modeling. In this paper, we propose MLICv2 and MLICv2+, enhanced successors that systematically address these limitations through improved transform design, advanced entropy modeling, and exploration of the potential of instance-specific optimization. For transform enhancement, we introduce a lightweight token mixing block inspired by the MetaFormer architecture, which effectively mitigates high-bit-rate performance degradation while maintaining computational efficiency. For entropy modeling improvements, we propose hyperprior-guided global correlation prediction to extract global context even in the initial slice of latent representation, complemented by a channel reweighting module that dynamically emphasizes informative channels. We further explore enhanced positional embedding and guided selective compression strategies for superior context modeling. Additionally, we apply the Stochastic Gumbel Annealing (SGA) to demonstrate the potential for further performance improvements through input-specific optimization. Extensive experiments demonstrate that MLICv2 and MLICv2+ achieve state-of-the-art results, reducing Bjøntegaard-Delta Rate by 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% on Kodak, Tecnick, and CLIC Pro Val datasets, respectively, compared to VTM-17.0 Intra.
@@ 1 Introduction
Image compression is a fundamental problem in multimedia communication, aiming to achieve compact representation while maintaining high reconstruction quality. Traditional codecs such as VVC Intra bross2021vvc rely on block-based hybrid coding frameworks refined over decades. However, as these hand-crafted methods approach performance saturation, learned image compression (LIC) has emerged as a new paradigm that leverages end-to-end optimization to outperform conventional codecs.
Recent LIC approaches balle2016end, theis2017lossy, balle2018variational, minnen2018joint, minnen2020channel, cheng2020learned, xie2021enhanced, he2021checkerboard, he2022elic, zou2022the, jin2025neural, zhao2023universal, tu2024reconstruction, zhu2022transformerbased, koyuncu2022contextformer, duan2023qarv, wang2022neural, guo2021causal, wu22block, yang2024perceptual, mishra2021multi, mishra2020wavelet, mishra2022deep2, mishra2022deep, park2023scalable, chen2021nic, liu2023learned, jiang2022mlic, jiang2023mlicpp, jiang2024llic typically follow an autoencoder-based framework optimized for the rate-distortion trade-off. An analysis transform encodes the image into latent representations, which are quantized and decoded by a synthesis transform. Entropy models then estimate the probability distribution of quantized latents to enable efficient coding. With advanced transform designs and entropy models, several LIC systems he2022elic, jiang2022mlic, jiang2023mlicpp, jiang2024llic, liu2023learned, li2024frequencyaware, lu2025learned, feng2025linear have achieved or even surpassed the performance of state-of-the-art traditional codecs such as VVC Intra bross2021vvc.
Entropy modeling is crucial to LIC performance as it minimizes conditional entropy by capturing informative priors. Hyperprior models balle2018variational introduce side information, while context models minnen2018image, he2021checkerboard capture spatial dependencies among neighboring elements. Building upon these foundations, Jiang et al. proposed the MLIC series jiang2022mlic, jiang2023mlicpp,
@@
Figure 1: BD-Rate and decoding time during inference on CLIC Professional Valid CLIC2020 with 2K resolution. The bubble size indicates the number of model parameters.
which partitions latents into channel-wise slices minnen2020channel to model both inter- and intra-slice local111“Local” denotes neighboring regions, specifically within the receptive field of a convolution or a windowed attention. and global222“Global” refers to the entire feature map. dependencies. Within each slice, checkerboard partitioning he2021checkerboard enables two-pass decoding, and attention modules model global correlations from previous slices. To reduce computational cost, MLIC++ jiang2023mlicpp replaces standard attention vas2017attention with linear attention shen2021efficient333“Linear” denotes complexity linear with spatial resolution., which scales more efficiently while maintaining global correlation modeling.
@@
Despite their strong performance, MLIC models still face several limitations that motivate this work. First, in terms of transform design, performance degrades at high bit-rates due to limited transform capacity, which restricts the preservation of latent information. Second, for entropy modeling, global correlation estimation relies solely on previously decoded slices, preventing the capture of global dependencies in early slices. Furthermore, all channels are treated equally without considering their varying informational importance, potentially reducing modeling accuracy.
@@
To systematically address these limitations, and as illustrated in Fig 2, we propose MLICv2, featuring enhanced transform and entropy modeling modules, and MLICv2+, which additionally incorporates iterative refinement techniques yang2020improving to demonstrate framework potential.
@@
For transform improvement, we design an efficient token mixing block to replace the residual blocks in MLIC++. Inspired by the MetaFormer architecture yu2022metaformer, this block employs depth-wise convolution for spatial mixing and lightweight gated interaction for point-wise operations, achieving better high-bit-rate performance with lower complexity.
@@
For entropy modeling, we propose MEMv2, a multi-reference entropy model that captures richer dependencies. We introduce hyperprior-guided global correlation prediction for initial slices by exploiting the spatial similarity between hyperprior and latent features. A two-stage context modeling strategy is also proposed: it first performs spatial aggregation and then applies adaptive channel-wise reweighting to capture feature importance variations. Further enhancements include 2D Rotary Positional Embedding and guided selective compression for improved modeling and coding efficiency.
@@
Finally, we employ Stochastic Gumbel Annealing (SGA) yang2020improving to iteratively refine latent representations and side information for each input, enabling input-specific rate-distortion optimization without increasing decoder complexity. This demonstrates the scalability and adaptability of our framework for future extensions.
@@
Our comprehensive approach leads to significant performance gains, as shown in Fig. 1. MLICv2 achieves BD-rate reductions of 16.54%, 21.61%, and 16.05% on the Kodak, Tecnick, and CLIC Professional Validation datasets, respectively, while MLICv2+ further improves these results to 20.46%, 24.35%, and 19.14% over VTM-17.0 Intra. Our contributions are summarized as follows:
Figure 2: Model evolution of the MLIC series.
@@
  • We introduce a MetaFormer-based transform architecture for LIC through an efficient token mixing block, achieving better high-bit-rate performance and lower complexity than previous residual designs.
@@
  • We enhance entropy modeling with three key innovations: (i) hyperprior-guided global correlation prediction for initial slices, (ii) the first two-stage context modeling strategy with spatial aggregation and adaptive channel reweighting, and (iii) advanced positional embedding and selective compression mechanisms.
  • We apply iterative refinement via SGA to demonstrate consistent rate-distortion improvements without increasing decoder complexity.
@@
  • Extensive experiments validate the effectiveness of MLICv2 and MLICv2+, achieving BD-rate reductions of 16.54%, 21.61%, 16.05% and 20.46%, 24.35%, 19.14% on the Kodak, Tecnick, and CLIC Professional Validation datasets, respectively.
@@ 2.1 Learned Image Compression
Autoencoder-based learned image compression balle2016end, theis2017lossy, minnen2018joint, cheng2020learned, he2022elic, jiang2022mlic, jiang2023mlicpp, minnen2020channel, jiang2024llic operates within the transform coding framework, where an input image $\boldsymbol{x}\sim p_{\boldsymbol{x}}$ is first mapped to a latent representation $\boldsymbol{y}=g_{a}(\boldsymbol{x};\phi)$ by an analysis transform $g_{a}$ with parameters $\phi$, then quantized as $\hat{\boldsymbol{y}}=\lceil\boldsymbol{y}\rfloor$ and finally reconstructed as $\hat{\boldsymbol{x}}=g_{s}(\hat{\boldsymbol{y}};\psi)$ using a synthesis transform $g_{s}$ with parameters $\psi$. A parametric entropy model $p(\cdot;\rho)$ with parameters $\rho$ is employed to estimate the latent distribution for efficient compression. During training, quantization is approximated by adding uniform noise $\boldsymbol{u}\sim\mathcal{U}(-0.5,0.5)$ or by using the straight-through estimator (STE) theis2017lossy. Following Ballé et al. balle2018variational, the training objective is formulated as:
$\mathcal{L}=\mathcal{R}(\hat{\boldsymbol{y}})+\lambda\,\mathcal{D}(\hat{\boldsymbol{x}},\boldsymbol{x})$}, (1)
[…]
Transform architectures have evolved significantly, incorporating Generalized Divisive Normalization (GDN) balle2015gdn, wavelets fu2024weconvene, mishra2020wavelet, attentions zhangresidual, chen2021nic, mishra2022deep2, multi-scale networks mishra2022deep, deep residual networks cheng2020learned, and transformer-based designs zhu2022transformerbased, zou2022the, lu2021tic. Mixed CNN-transformer liu2023learned and Mamba zeng2025mambaic, gumamba, qin2024mambavc architectures have also been explored for capturing both local and non-local interactions. Entropy estimation has progressed from simple univariate Gaussian models balle2018variational to sophisticated approaches including mean-scale Gaussian models minnen2018joint, asymmetric Gaussian distributions cui2021asym, Gaussian mixture models cheng2020learned, and generalized Gaussian models zhang2025generalized, significantly improving compression efficiency.
@@ 2.2 Conditional Entropy Models
Ballé et al. balle2018variational introduce hyperprior networks using hyper analysis $h_{a}$ and hyper synthesis $h_{s}$ to improve entropy estimation through quantized side information $\hat{\boldsymbol{z}}$ extracted from latent representations.
Local context models have evolved from serial methods, such as PixelCNN van2016conditional, minnen2018joint, to more efficient parallel approaches, including zigzag scanning li2020efficient and checkerboard partitioning he2022elic, which focus on intra-slice dependencies by treating the latent representation as a slice. Minnen et al. minnen2020channel introduced channel-wise autoregressive coding to efficiently model dependencies along the channel dimension and achieve faster decoding.
Recent studies further extend context modeling to capture long-range dependencies qian2020learning, guo2021causal, kim2022joint, jiang2022mlic, jiang2023mlicpp. For instance, Qian et al. qian2020learning, qian2022entroformer introduce bi-directional attention to model global correlations, while Guo et al. guo2021causal and Kim et al. kim2022joint transmit global priors directly to the decoder for enhanced reconstruction.
Recently, Jiang et al. jiang2022mlic propose MLIC and MLIC+, dividing global contexts into inter-slice and intra-slice components. The latent representation is partitioned into multiple slices for autoregressive conditional coding, employing checkerboard attention for intra-slice contexts and convolutional layers for inter-slice contexts. However, the quadratic complexity of vanilla attention vas2017attention limits scalability for high-resolution images, which is addressed by MLIC++ jiang2023mlicpp through linear attention mechanisms shen2021efficient.
Figure 3: The overall architecture of MLICv2/MLICv2+. $g_{a}$ is the analysis transform, $g_{s}$ is the synthesis transform, $h_{a}$ is the hyper analysis, and $h_{s}$ is the hyper synthesis. $g_{ep}$ is the entropy parameter module. $f_{s}$ is the selective compression predictive module. $g_{lrp}$ is the latent residual prediction module minnen2020channel. AE and AD are arithmetic encoding and decoding. $\uparrow$ denotes upsampling and $\downarrow$ denotes downsampling. $\boldsymbol{x}$} is the input image and $\hat{\boldsymbol{x}}$ is the reconstructed image. $\boldsymbol{y}$ is the latent representation and $\hat{\boldsymbol{y}}$ is the quantized latent representation. $\hat{\boldsymbol{y}}^{i}$ is the $i$-th slice of $\hat{\boldsymbol{y}}$. $\boldsymbol{\mu}^{i},\boldsymbol{\sigma}^{i}$ are the mean and scale of $\hat{\boldsymbol{y}}^{i}$. $\boldsymbol{s}^{i}$ is the predicted selective coding map. $\boldsymbol{r}^{i}$ is the predicted residual. $\hat{\boldsymbol{z}}$ is the side information. $M,N$ are the channel numbers, which are $320,192$ in MLICv2/MLICv2+. “Refine” is employed in MLICv2+.
@@ 3.1 Background: MLIC++
Our proposed MLICv2 and MLICv2+ are built upon the foundation of MLIC++ jiang2023mlicpp. MLIC++ adopts a simplified variant of Cheng’20’s transform cheng2020learned, where attention modules are removed to reduce computational complexity. For entropy modeling, MLIC++ introduces a linear-complexity multi-reference entropy model (MEM++) that effectively captures hyperprior, intra-slice, and inter-slice dependencies.
The hyperprior information $\hat{\boldsymbol{H}}$ is derived from side information $\hat{\boldsymbol{z}}$. Following the approach of Minnen et al. minnen2020channel, the latent representation $\hat{\boldsymbol{y}}$ is partitioned into $L+1$ slices: $\hat{\boldsymbol{y}}^{0},\hat{\boldsymbol{y}}^{1},\ldots,\hat{\boldsymbol{y}}^{L}$}. Each slice $\hat{\boldsymbol{y}}^{i}$ is further divided into anchor $\hat{\boldsymbol{y}}^{i}_{ac}$ and non-anchor $\hat{\boldsymbol{y}}^{i}_{na}$ components using checkerboard partitioning he2021checkerboard.
The inter-slice local context module $g_{\ell}^{inter}$ extracts inter-slice local context $\hat{\boldsymbol{C}}^{<i}_{\ell}$ from preceding slices $\hat{\boldsymbol{y}}^{<i}$. To model inter-slice global context $\hat{\boldsymbol{C}}^{<i}_{g}$, MLIC++ employs linear attention shen2021efficient to achieve an optimal balance between performance and computational complexity. The intra-slice local context $\hat{\boldsymbol{C}}^{i}_{\ell}$ is extracted from the anchor partition $\hat{\boldsymbol{y}}^{i}_{ac}$ using overlapped window-based checkerboard attention $g_{\ell}^{intra}$.
@@
The estimated means $\boldsymbol{\mu}^{i}_{ac}$, $\boldsymbol{\mu}^{i}_{na}$ and scales $\boldsymbol{\sigma}^{i}_{ac}$, $\boldsymbol{\sigma}^{i}_{na}$ are extracted from the contextual information using the entropy parameter module $g_{ep}$. MLIC++ incorporates mixed quantization minnen2020channel and the latent residual prediction module $g_{lrp}$ minnen2020channel. In mixed quantization, zero-centered straight-through estimation (STE) theis2017lossy is applied for distortion computation, while additive uniform noise (AUN) balle2016end is used for entropy estimation during training. The module $g_{lrp}$ predicts quantization residuals based on decoded slices and hyperprior information: